OpenStudy (anonymous):

Eliminate the parameter t. Find a rectangular equation for the plane curve defined by the parametric equations. x = 6 cos t, y = 6 sin t; 0 ≤ t ≤ 2π A. x^2 - y^2 = 6; -6 ≤ x ≤ 6 B. x^2 - y^2 = 36; -6 ≤ x ≤ 6 C. x^2 + y^2 = 36; -6 ≤ x ≤ 6

OpenStudy (anonymous):

since \(\cos^2(t)+\sin^2(t)=1\) you have \(x^2+y^2=36\)

OpenStudy (anonymous):

parametric equation of circle:)

OpenStudy (anonymous):

so it would be c?

OpenStudy (anonymous):

yup:)

OpenStudy (anonymous):

thank you